If xe xy - y = sin 2 x , then find dy/dx at x = 0. Please answer this question asap with proper steps so that a grade 11 student can understand. thank u.

If xexy - y = sin2x , then find dy/dx at x = 0. Please answer this question asap with proper steps so that a grade 11 student can understand. thank u.
 
 

Grade:11

1 Answers

Arun
4 years ago
Given eqn.  xe^(xy) – y = sin^2x,
Differentiate y  w.r.t.   x as:

(1)e^(xy) + xe^(xy)*[1.y + xdy/dx] – dy/dx = sin2x
collect dy/dx,

put x = 0:
  dy/dx =  1 is the answer